DATA AND ANALYSIS OF SPIN-ORBIT COUPLED A^{1}Σ_{u}^{+} AND b^{3}Π_{u} STATES OF Cs_{2}

\underline{ANDREY V. STOLYAROV}^{a}, Department of Chemistry, Moscow State University, GSP-2 Leninskie gory 1/3, Moscow 119992, Russia; THOMAS H. BERGEMAN, Department of Physics and Astronomy, State University of New York, Stony Brook, New York 11794-3800.

Following work with various collaborators on analogous states of Na_{2}, NaRb, KRb, NaCs, KCs, RbCs and Rb_{2}, we have recently completed a study of the A^{1}Σ_{u}^{+} and b^{3}Π_{u} states of Cs_{2}, based on experimental data obtained in Orsay, Temple University, Innsbruck and Riga. Potentials and spin-orbit coupling functions were fit to the data using both a coupled-channel approach and the discrete variable representation method. As with other species, the goal has been to accurately characterize the energy level structure to facilitate experiments in which cold ground state molecules are produced from cold atoms or Feshbach resonance dimers by stepwise excitation and decay, and also to facilitate excitations to higher lying states. For Cs_{2}, second-order spin-orbit shifts become significant, and since these are typically difficult to obtain experimentally, ab initio calculations provided estimates. We will display the excellent quality of the fits to the experimental data, including a modeling of the b^{3}Π_{1u}-^{3}Π_{0u}^{+} fine structure interval and Λ-doubling effect in the b^{3}Π_{0u}^{±} state.

aSupport by RFBR is gratefully acknowledged